OpenStudy (anonymous):

log(6)36x use properties of logarithms to expand the logarithmic expression as much as possible.

OpenStudy (anonymous):

another type to try to remember!

OpenStudy (anonymous):

Re-write it as a power.

OpenStudy (anonymous):

36x^6 ??

OpenStudy (anonymous):

wait, scratch that

OpenStudy (anonymous):

\[\large \log_b(a) = k \iff b^k = a\]

OpenStudy (anonymous):

So here we have: \[\large \log_6(36x) \iff \]

OpenStudy (anonymous):

6^x=36

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Err sorry. Let's let k = that log. \[\large \log_6(36x) = k \iff\]

OpenStudy (anonymous):

Bah! wait

OpenStudy (anonymous):

but where does k jump in from?

OpenStudy (anonymous):

I made it up. But it's not going to be helpful

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Sorry. I think they want you to see that 36 is a power of the base.

OpenStudy (anonymous):

6^2=36 ???

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

\[\log_b\left(b^2\cdot x\right) = \log_b\left(b^2\right) + log_b(x) \]

OpenStudy (anonymous):

Then you can simplify the first log there

OpenStudy (anonymous):

\[log_b(b^2) = 2log_b(b) = 2\]

OpenStudy (anonymous):

memorizing all these formulas is what gets me... maybe I need a cheat sheet!

OpenStudy (anonymous):

it all looks like gibberish to me :(

OpenStudy (anonymous):

The cheat sheet is this: \[log_b(a) = k \iff b^k = a\]

OpenStudy (anonymous):

From that rule, (and all the laws of exponents) you can find all the laws of logs.

OpenStudy (anonymous):

But it mostly takes practice

OpenStudy (anonymous):

my brain is very linguistic vs numeric

OpenStudy (anonymous):

so is my answer then 2? it seems too simple

OpenStudy (anonymous):

Symbols, not numbers.. my explanations have no numbers ;p

OpenStudy (anonymous):

It's not 2. It's 2 + log_6(x)

OpenStudy (anonymous):

But hold on, cause you'll need to know this later.

OpenStudy (anonymous):

Try this on your linguistic brain.. "The log (base b) of some value a, is the exponent you must use on b in order to get a."

OpenStudy (anonymous):

So if I ask you , what's the log (base 6) of 36 ?

OpenStudy (anonymous):

one second if you don't mind... got an important call. i'll b back!!

OpenStudy (anonymous):

sry... lost my train of thought... business call.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

okay, so I gravitate back to log(2)

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

No, try this I want to know what the log (base 5) of 125 is.

OpenStudy (anonymous):

geesh, i dunno. you asked a question up there. this stuff is beyond me it seems!

OpenStudy (anonymous):

ok, lemme see

OpenStudy (anonymous):

Now remember what I said about what the log is.. "The log (base b) of some value a, is the exponent you must use on b in order to get a."

OpenStudy (anonymous):

log(5)125=3

OpenStudy (anonymous):

Because?

OpenStudy (anonymous):

5 to the power of 3 = 125

OpenStudy (anonymous):

Right. 3 is the exponent you must use on 5 (the base) to get 'a' (125)

OpenStudy (anonymous):

So now lets go back to our example here.. Lets just imagine that there exists some value (we'll call it k) for which \[\large\log_6(36x) = k \]

OpenStudy (anonymous):

Now, if k is the log of that inside part. that means what?

OpenStudy (anonymous):

log(6)36x=2

OpenStudy (anonymous):

No, it doesn't. It equals k

OpenStudy (anonymous):

ohhhhh, ok

OpenStudy (anonymous):

It means that \[\large6^k = 36x\] right?

OpenStudy (anonymous):

ok, I see that.

OpenStudy (anonymous):

Ok, but we also know that \(36 = 6^2\)

OpenStudy (anonymous):

So lets write it that way.. \[\implies 6^k = 6^2x\]

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

yes, i see the logic there.

OpenStudy (anonymous):

Ok, so now lets pretend we know what k is, but we're still wondering about x

OpenStudy (anonymous):

We can 'solve' for x by dividing both sides by \(6^2\)

OpenStudy (anonymous):

\[\frac{6^k}{6^2} = x\]

OpenStudy (anonymous):

And how do you simplify \(\Large\frac{6^k}{6^2} = x\)

OpenStudy (anonymous):

'When you divide powers of the same base you subtract the exponent of the denominator from the exponent of the numerator'

OpenStudy (anonymous):

\[\large\implies 6^{k-2} = x\] And now we take the log base 6 of both sides.. \[\implies k-2 = \log_6(x)\] Solve for k \[k = 2+\log_6(x)\] And since we said originally that \(\large k = \log_6(36x)\) we have: \[\large \implies \log_6(36x) = 2 + log_6(x)\]

OpenStudy (anonymous):

But really this is a very meandery way of simplifying. Once you have more practice with logs (rather than getting answers) you will start to see easier short cuts: \[\qquad \qquad \log_6(36x)\]\[=\log_6(36) + \log_6(x)\]\[=\log_6(6^2) + log_6(x)\]\[=2 + log_6(x)\]

OpenStudy (anonymous):

whoa... it's just overwhelming.

OpenStudy (anonymous):

I can't imagine retaking the course again after 16 weeks of brain numbing algebra!

OpenStudy (anonymous):

Algebra is good exercise

OpenStudy (anonymous):

so, thank you for the help! I am so greatful to all the good helpers here

OpenStudy (anonymous):

you're right, exercise for the brain, I felt that way in the beginning but fried with this advanced stuff.

OpenStudy (anonymous):

\[\qquad \qquad \log_6(36x)\](The log of a product is the sum of the log of each factor)\[= \log_6(36) + \log_6(x)\]\[=\log_6(6^2) + \log_6(x)\](The log of a power is the product of the exponent and the base of the power)\[=2\log_6(6) + \log_6(x)\](The log(base b) of b is 1)\[=2(1) + \log_6(x) = 2 + \log_6(x)\]

OpenStudy (anonymous):

i'm going to print that out :)